# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v2

@@ 4.1. Containers in Multi-Cloud: Roles and Strategies (RQ1)
Multifaceted Role of Containers in Multi-Cloud Environment: Containers fulfill diverse roles within multi-cloud environment, such as application deployment, service hosting, and resource management. They facilitate deployment of applications across a broad range of systems, from web apps to IoT applications. Furthermore, containers contribute to enhancing security in multi-cloud operations. Implications: Based on these findings, researchers have opportunities to investigate further areas, such as potential security vulnerabilities, management complexities, and performance overheads. Similarly, practitioners can use this understanding to inform efficient and effective deployment strategies, leading to more robust multi-cloud implementations.
@@
Container Technology Features and their Significance: The unique features of container technology, such as lightweight virtualization, compatibility with orchestration tools, and application in edge computing, offer significant advantages. Implications: Researchers can explore these features further, including how characteristics like lightweight virtualization influence container performance and effectiveness in multi-cloud operations. Practitioners can leverage this understanding for better utilization of container technology and more efficient multi-cloud deployments.
@@
Implementation Strategies and their Real-World Implications: The strategic implementation of containers significantly affects their functionality in multi-cloud environment. Strategies related to deployment, orchestration, security, performance optimization, cloud service management, and data storage can have significant implications. Implications: Researchers may explore these strategies further, including investigations into their real-world implications, benefits, and limitations. Practitioners can refine their understanding of these strategies to achieve more efficient implementations and operations.
@@
Containerization’s Impact on IoT and Application Development: The incorporation of containers into IoT and app development has transform these fields. Containers offer improved software unit encapsulation, portability, and efficiency, which are particularly important in the distributed and resource-constrained nature of IoT. Implications: Researchers can further explore potential challenges in detail such as network complexity, security, and the small footprint requirement of IoT devices. For practitioners, understanding the impact of containerization can lead to the development of more robust, efficient, and secure IoT and app solutions.
4.2. Patterns and Strategies (RQ2)
@@
Diverse Architectural Patterns and Communication Strategies: Our study found that Microservice Architecture and Service-Oriented Architecture (SOA) are the most common used architectural patterns for designing container-based applications in multi-cloud environment. These results align with the existing literature that emphasizes the use of MSA and SOA because of their support for modular development, scalability, and efficient resource use (Balalaie et al., 2016; Dragoni et al., 2017). The other patterns frequently reported are Service Mesh and Service Chaining, which are employed as communication and networking patterns, likely due to their ability to manage complex, distributed service-to-service communications in a scalable way (Zhamak, 2018). The frequent use of black-Green Deployment and Object Store Service deployment patterns has also been identified in this study, which may stem from their ability to minimize downtime and ensure high availability—essential factors in a cloud-based setting (fowler2010blackgreen). Implications: Practitioners can enhance the design, scalability, and efficiency of their applications in a multi-cloud environment by understanding and applying these architectural and communication patterns. Researchers might explore how these patterns evolve with the emergence of new technologies and standards.
@@
Importance of Container Management and Multi-Cloud Strategies: Our findings report the significance of robust container management and multi-cloud strategies. Techniques like the Linux Container (LXC) project, Docker Container Images, and Lightweight Virtualization are frequently used, emphasizing the industry trend towards containerization due to its benefits in efficiency, scalability, and platform independence (Vaughn, 2020). Furthermore, the popularity of multi-cloud strategies like Hybrid/Multi-cloud Approach and Multi-cloud Load Balancing confirms the growing industry move towards using multiple cloud providers to enhance redundancy, flexibility, and avoid vendor lock-in (Gillam et al., 2013). Implications: Practitioners should focus on gaining expertise in container technologies and multi-cloud strategies to take advantage of their numerous benefits. Meanwhile, researchers might examine the challenges and best practices associated with the deployment and management of container technologies across multiple cloud platforms.
Security, Resilience, and Migration Patterns and Strategies: Our study also identifies patterns and strategies related to security, resilience, and efficient migration strategies in a multi-cloud context. Strategies such as File Encryption, Attribute-Based Encryption, and Secure Data sharing are often employed to ensure data security, an area of growing concern with the increase in cloud-based applications (Svantesson and Clarke, 2010). Moreover, resilience strategies such as Fault-tolerance with Redundant Engines and RAFT Consensus Algorithm are essential to ensure system reliability in distributed, multi-cloud environment (Ongaro and Ousterhout, 2014).
@@
Migration between different cloud environments, facilitated by strategies like Service-oriented Migration and Migration to a Virtualized Container, is another vital aspect, given the growth in cloud services and the need for interoperability (Jamshidi et al., 2013). Migrating applications and data between cloud environments is an important consideration in a multi-cloud scenario (Jamshidi et al., 2013). Strategies such as image-based migration and migration to virtualized containers, like Docker, are commonly used (Vaughn, 2020). These techniques facilitate portability and provide consistent environments across different cloud platforms. Additionally, they have significant implications for scaling and managing applications across different cloud service providers. Implications: These findings suggest that security, resilience, and efficient migration should be key areas of focus for practitioners operating in multi-cloud environment. Researchers may also look into novel techniques for enhancing cloud security, improving system resilience, and easing the process of cloud migration.
4.3. Quality Attributes and Tactics (RQ3)
@@
Performance Optimization in Multi-Cloud Environment: Our study identifies the role of machine learning, container technology, and location-aware service brokering in enhancing the performance and efficiency of containerized applications within multi-cloud environment. Hashem et al. previously suggested similar efficiency gains from machine learning and container technology, validating the interpretive value of our results (Hashem et al., 2016). Specifically, the application of machine learning aligns with a broader trend towards intelligent cloud resource management, a domain that Buyya et al. identified as significant in cloud computing research (Buyya et al., 2020). Meanwhile, location-aware service brokering emerges as a novel and impactful strategy, stressing the role of geographical considerations in further optimizing containerized applications’ performance. Implications: For researchers, the results of our study offer opportunities to explore the integration of AI techniques with container-based cloud architectures further, especially the application of location-aware service brokering in multi-cloud environment. For practitioners, the results indicate that a adaptable approach involving AI, container technology, and geographical considerations can lead to significant performance enhancements in multi-cloud environment.
@@
Security Measures in Multi-Cloud Deployments: The results of our study indicates the multi-faceted approach to ensuring security in multi-cloud deployments. It highlights the combination of encryption, network security protocols, machine learning, and user-specified security measures, while also encouraging deployment across different cloud providers. This finding complements Singh et al.’s argument for a comprehensive approach to cloud security, and aligns with the recent concept of distributed cloud security (Singh et al., 2016). Our results also report the importance of user-specified security measures, which are in agreement with the user-centric security model proposed by Shin and Dong-Hee (Shin, 2013). Implications: From a research perspective, the interact between different security measures and their effect on overall system security in multi-cloud environment warrants further investigation. For practitioners, these results point out the importance of employing a comprehensive security approach in multi-cloud deployments, including encryption, network security, machine learning, and user-specific security measures.
@@
Ensuring Compatibility Across Different Cloud Providers: The results of our study suggests strategies to ensure compatibility in multi-cloud environment through standardization of interfaces, componentization, and interoperability approaches. Emphasizing lightweight communication protocols and interoperability between clouds, the results related to compatibility mirrors Mell and Grance’s principles of service-oriented architecture, reinforcing the need for effective communication across different cloud platforms (Mell et al., 2011). Petcu et al. have also previously stressed the importance of standardization and componentization for ensuring compatibility (Petcu et al., 2018). Implications: From a research standpoint, the dynamic between lightweight communication protocols, interface standardization, and overall system compatibility in multi-cloud deployments offers fertile ground for future work. Practitioners, meanwhile, can adopt these approaches to ensure seamless communication and interoperability across different cloud providers.
@@
4.4. Automation Challenges and Solution Framework (RQ4 - RQ5)
@@
Multi-dimensional Challenges: The Automation Challenge-Solution Framework for Containerized Applications in multi-cloud environment suggests that multi-cloud management is a complex task encompassing several dimensions. These include multi-cloud automation, deployment and scaling, resource management, data and application migration, testing and bench marking, standardization and interoperability, application and service management, and runtime and service discovery. This indicates that addressing one challenge may potentially impact the other areas, necessitating a holistic approach to tackle these issues. The multi-dimensionality of the challenges implies the need for an integrated approach that considers the interdependencies among these dimensions. It emphasizes the need for solutions that cater to this complexity rather than address individual challenges in isolation. Implications: For researchers, these findings underscore the need to develop multi-faceted, holistic solutions that can address the various interconnected challenges in managing multi-cloud environment. For practitioners, this implies the importance of strategic planning and implementing solutions that address the multiple facets of multi-cloud management while considering their potential impact on one another.
@@
Emergence of AI/ML Solutions: The framework highlights the increasing utilization of AI and ML in providing intelligent solutions to manage containerized applications in multi-cloud environment. This suggests the growing maturity and applicability of these technologies in complex scenarios, a significant step towards achieving efficient multi-cloud management. Previous studies (e.g., (de Laat et al., 2020), (Chen et al., 2021), (Gill et al., 2022)) have recognized the potential of AI and ML in resolving complex cloud management issues, which is in line with the findings from the framework. This further validates the significant role AI and ML have started to play in this domain. The increasing adoption of AI/ML in multi-cloud environment signifies an essential evolution in the field. This stresses the need for further research and development in these technologies to exploit their full potential in resolving multi-cloud management complexities. Implications: The growing use of AI/ML solutions implies that researchers need to focus on improving these technologies and adapting them to specific multi-cloud management challenges. For practitioners, this points towards the increasing necessity to invest in AI/ML capabilities and integrate them into their multi-cloud management strategies.
@@
Importance of Standardization and Interoperability: The framework underscores the crucial role of standardization and interoperability in multi-cloud environment, facilitating seamless integration and management across different cloud platforms. Recent literature validates the importance of standardization and interoperability in multi-cloud environment, emphasizing their role in enhancing the efficiency of cloud services (Alonso et al., 2023b). This agrees with our results, further confirming the critical need for these factors in the domain. The need for standardization and interoperability reiterates the necessity for a unified framework or protocol across different cloud environments. It suggests that attention to these aspects can lead to more seamless, efficient, and secure management of multi-cloud services. Implications: For researchers, the results stress the need to investigate methods to achieve better standardization and interoperability in multi-cloud environment. For practitioners, the findings suggest the need to adopt standards and focus on interoperability while selecting and implementing cloud services, ensuring more efficient management across different platforms.
4.5. Security Challenges and Solutions Framework (RQ4 - RQ5)
@@
Emphasize on Container-Specific Security Mechanisms: Security solutions specifically designed for containerized applications are critical to managing threats in multi-cloud environment. The selected studies suggest a variety of methods, including secure container orchestration, container isolation, and advanced deployment strategies that make use of deep reinforcement learning. This aligns with the existing literature (e.g., (Zhong et al., 2022), (Li et al., 2023) (Combe et al., 2016)) which highlights the unique security considerations brought forth by containerization. Containers, being lighter than traditional virtual machines and offering process-level isolation, have revolutionized the way applications are packaged and run. However, they also introduce new security challenges that need to be addressed by using container-specific security mechanisms. Containerization, while a powerful tool for application deployment, necessitates its own suite of security strategies. Implementing these will be key to leveraging the benefits of containerization in a secure manner. Implications: For researchers, this highlights an area of potential study: the development of advanced container-specific security strategies, including those that use advanced techniques such as machine learning. For practitioners, this suggests the need to be well-versed in the specifics of container security. Leveraging features of containerization platforms, such as Docker’s isolation features, or incorporating tools specifically designed for container security will be crucial.
@@
Implementing Multi-Layered Security and Compliance Measures: Implementing multi-layered security measures that encompass data protection, access control, and communication security, along with compliance with legal and regulatory requirements, is of utmost importance. Solutions range from data encryption, role-based access control, secure container orchestration, adherence to regulations like GDPR (Regulation, 2016). This is consistent with current cybersecurity frameworks, such as the NIST cybersecurity framework (of Standards and Technology, 2018), which emphasize a multi-layered approach to security. Furthermore, several studies point to the criticality of data security, user access control, and secure communication in cloud environments (Austin, 2018) (Singh et al., 2016). In a multi-cloud, containerized environment, it is not enough to secure data; access and communication channels must also be secured, and all actions must comply with legal and regulatory requirements. Implications: For practitioners is to design and implement a comprehensive security strategy that encompasses multiple layers of protection. Compliance with legal and industry standards should be an integral part of this strategy. For researchers, this underlines the need for studies that address the integration of various layers of security and the development of comprehensive security frameworks.
@@
Ensuring Security through Standardization and Interoperability: Ensuring security through standardization and interoperability is highlighted as a critical step. Secure and standardized interfaces, such as OpenStack APIs, can be used to enhance security and performance across multiple cloud providers. This finding is aligned with existing literature, which emphasizes the role of standardization in achieving security and interoperability in multi-cloud environment (Buyya et al., 2010) (Petcu, 2011). By implementing standardized interfaces and practices, organizations can ensure a level of interoperability and security across cloud environments. Implications: For practitioners, this means choosing platforms and tools that support standardized interfaces and practices, or working to implement those standards within their own environments. Researchers, on the other hand, could focus on the development of new standards or the improvement of existing ones to better address security challenges in multi-cloud, containerized environments.
4.6. Deployment Challenges and Solutions Framework (RQ4 - RQ5)

### v3

@@ 4.1 Containers in Multi-Cloud: Roles and Strategies (RQ1)
Multifaceted Role of Containers in Multi-Cloud Environment: In a multi-cloud environment, containers serve multiple purposes, such as managing resources, hosting services, and deploying applications. They simplify the deployment of apps across various platforms, including web applications and Internet of Things (IoT) applications. Additionally, containers enhance security in multi-cloud operations. These roles are shaped by the inherent heterogeneity of multi-cloud settings, where containers are used to ensure portability, manage provider-specific constraints, and support workload migration across isolated cloud infrastructures. Implications: These findings open up opportunities for researchers to explore additional areas, including security vulnerabilities, management challenges, and performance overhead. Likewise, practitioners can apply these insights to design more efficient and effective deployment strategies, ultimately contributing to stronger and more reliable multi-cloud implementations.
@@
Container Technology Features and their Significance: Container technology offers several benefits due to its unique features, such as support for edge computing, compatibility with orchestration tools, and lightweight virtualization. In multi-cloud environments, these features enable containers to act as a unifying abstraction layer across diverse infrastructure stacks, allowing consistent deployment behavior despite differing cloud-specific configurations. Implications: Researchers can explore these features further, including how characteristics like lightweight virtualization influence container performance and effectiveness in multi-cloud operations. Practitioners can use this understanding for better utilization of container technology and more efficient multi-cloud deployments.
@@
Implementation Strategies and their Real-World Implications: The strategic implementation of containers significantly affects their functionality in multi-cloud environment. Strategies related to deployment, orchestration, security, performance optimization, cloud service management, and data storage can have significant implications. The reviewed strategies demonstrate how container management must adapt to cloud-specific APIs, variable service guarantees, and cross-cloud policy enforcement, revealing how the interplay between containerization and multi-cloud governance introduces distinctive design trade-offs. Implications: Researchers may explore these strategies further, including investigations into their real-world implications, benefits, and limitations. Practitioners can refine their understanding of these strategies to achieve more efficient implementations and operations.
@@
Containerization’s Impact on IoT and Application Development: The incorporation of containers into IoT and app development has transform these fields. Containers offer improved software unit encapsulation, portability, and efficiency, which are particularly important in the distributed and resource-constrained nature of IoT. In multi-cloud scenarios, this impact is increase by the need to coordinate deployments across heterogeneous edge-cloud setups and maintain interoperability between varied cloud-native services and orchestration frameworks. Implications: Researchers can further explore potential challenges in detail such as network complexity, security, and the small footprint requirement of IoT devices. For practitioners, understanding the impact of containerization can lead to the development of more robust, efficient, and secure IoT and app solutions.
4.2 Patterns and Strategies (RQ2)
@@
Diverse Architectural Patterns and Communication Strategies: Our study found that Microservice Architecture and Service-Oriented Architecture (SOA) are the most common used architectural patterns for designing container-based applications in multi-cloud environment. These results align with the existing literature that emphasizes the use of MSA and SOA because of their support for modular development, scalability, and efficient resource use [149, 150]. The other patterns frequently reported are Service Mesh and Service Chaining, which are employed as communication and networking patterns, likely due to their ability to manage complex, distributed service-to-service communications in a scalable way [151]. The frequent use of black-Green Deployment and Object Store Service deployment patterns has also been identified in this study, which may stem from their ability to minimize downtime and ensure high availability—essential factors in a cloud-based setting [152]. Notably, these architectural and communication patterns are adapted in the multi-cloud setting to support distributed orchestration, cross-cloud resilience, and integration across heterogeneous platforms—attributes that are less emphasized in single-cloud deployments. For instance, service mesh implementations in multi-clouds enable secure communication across isolated cloud boundaries, and decentralized orchestration patterns accommodate independently governed cloud nodes. Implications: Practitioners can enhance the design, scalability, and efficiency of their applications in a multi-cloud environment by understanding and applying these architectural and communication patterns. Researchers might explore how these patterns evolve with the emergence of new technologies and standards.
@@
Importance of Container Management and Multi-Cloud Strategies: Our findings report the significance of robust container management and multi-cloud strategies. Techniques like the Linux Container (LXC) project, Docker Container Images, and Lightweight Virtualization are frequently used, emphasizing the industry trend towards containerization due to its benefits in efficiency, scalability, and platform independence [153]. Furthermore, the popularity of multi-cloud strategies like Hybrid/Multi-cloud Approach and Multi-cloud Load Balancing confirms the growing industry move towards using multiple cloud providers to enhance redundancy, flexibility, and avoid vendor lock-in [154]. In particular, the reviewed studies illustrate how container technologies are adapted to meet challenges specific to multi-cloud deployments, such as orchestrating containers across provider-specific APIs, aligning container runtime environments across clouds, and ensuring policy compliance in cross-provider setups. These represent key peculiarities introduced by multi-cloud environments into the containerization landscape. Implications: Practitioners should focus on gaining expertise in container technologies and multi-cloud strategies to take advantage of their numerous benefits. Meanwhile, researchers might examine the challenges and best practices associated with the deployment and management of container technologies across multiple cloud platforms.
@@
Security, Resilience, and Migration Patterns and Strategies: Our study also identifies patterns and strategies related to security, resilience, and efficient migration strategies in a multi-cloud context. Strategies such as File Encryption, Attribute-Based Encryption, and Secure Data sharing are often employed to ensure data security, an area of growing concern with the increase in cloud-based applications [155]. Moreover, resilience strategies such as Fault-tolerance with Redundant Engines and RAFT Consensus Algorithm are essential to ensure system reliability in distributed, multi-cloud environment [156]. Migration between different cloud environments, facilitated by strategies like Service-oriented Migration and Migration to a Virtualized Container, is another vital aspect, given the growth in cloud services and the need for interoperability [157]. Migrating applications and data between cloud environments is an important consideration in a multi-cloud scenario [157]. Strategies such as image-based migration and migration to virtualized containers, like Docker, are commonly used [153]. These techniques facilitate portability and provide consistent environments across different cloud platforms. Additionally, they have significant implications for scaling and managing applications across different cloud service providers. Unlike traditional container security or resilience strategies in single-cloud contexts, the reviewed studies show that in multi-cloud scenarios, these approaches must handle cross-cloud identity management, consistency in fault detection and recovery across isolated infrastructures, and security rule federation across providers. Similarly, migration patterns uniquely address compatibility mismatches and compliance constraints introduced by the heterogeneity of cloud APIs and services. Implications: These findings suggest that security, resilience, and efficient migration should be key areas of focus for practitioners operating in multi-cloud environment. Researchers may also look into novel techniques for enhancing cloud security, improving system resilience, and easing the process of cloud migration.
4.3 Quality Attributes and Tactics (RQ3)
@@
Performance Optimization in Multi-Cloud Environment: Our study identifies the role of machine learning, container technology, and location-aware service brokering in enhancing the performance and efficiency of containerized applications within multi-cloud environment. Hashem et al. previously suggested similar efficiency gains from machine learning and container technology, validating the interpretive value of our results [158]. Specifically, the application of machine learning aligns with a broader trend towards intelligent cloud resource management, a domain that Buyya et al. identified as significant in cloud computing research [159]. Meanwhile, location-aware service brokering emerges as a novel and impactful strategy, stressing the role of geographical considerations in further optimizing containerized applications’ performance. Importantly, the reviewed studies emphasize that performance tactics must accommodate multi-cloud-specific challenges such as distributed latency, resource fragmentation across providers, and dynamic SLA enforcement, which are less relevant in single-cloud settings. Implications: For researchers, the results of our study offer opportunities to explore the integration of AI techniques with container-based cloud architectures further, especially the application of location-aware service brokering in multi-cloud environment. For practitioners, the results indicate that a adaptable approach involving AI, container technology, and geographical considerations can lead to significant performance enhancements in multi-cloud environment.
@@
Security Measures in Multi-Cloud Deployments: The results of our study indicates the multi-faceted approach to ensuring security in multi-cloud deployments. It highlights the combination of encryption, network security protocols, machine learning, and user-specified security measures, while also encouraging deployment across different cloud providers. This finding complements Singh et al.’s argument for a comprehensive approach to cloud security, and aligns with the recent concept of distributed cloud security [160]. Our results also report the importance of user-specified security measures, which are in agreement with the user-centric security model proposed by Shin and Dong-Hee [161]. Notably, multi-cloud containerization introduces additional layers of complexity in managing security policies across heterogeneous platforms and enforcing access control in federated settings. Tactics such as decentralized IAM logging and policy-aware container placement directly respond to these unique multi-cloud constraints. Implications: From a research perspective, the interact between different security measures and their effect on overall system security in multi-cloud environment warrants further investigation. For practitioners, these results point out the importance of employing a comprehensive security approach in multi-cloud deployments, including encryption, network security, machine learning, and user-specific security measures.
@@
Ensuring Compatibility Across Different Cloud Providers: The results of our study suggests strategies to ensure compatibility in multi-cloud environment through standardization of interfaces, componentization, and interoperability approaches. Emphasizing lightweight communication protocols and interoperability between clouds, the results related to compatibility mirrors Mell and Grance’s principles of service-oriented architecture, reinforcing the need for effective communication across different cloud platforms [162]. Petcu et al. have also previously stressed the importance of standardization and componentization for ensuring compatibility [163]. Our findings highlight that achieving compatibility in a containerized multi-cloud context often involves abstracting away provider-specific interfaces through standardized APIs and decoupling infrastructure logic via containerization. These are adaptations not typically required in homogeneous cloud environments. Implications: From a research standpoint, the dynamic between lightweight communication protocols, interface standardization, and overall system compatibility in multi-cloud deployments offers fertile ground for future work. Practitioners, meanwhile, can adopt these approaches to ensure seamless communication and interoperability across different cloud providers.
@@
4.4 Automation Challenges and Solution Framework (RQ4 - RQ5)
@@
Multi-dimensional Challenges: The Automation Challenge-Solution Framework for Containerized Applications in multi-cloud environment suggests that multi-cloud management is a complex task encompassing several dimensions. These include multi-cloud automation, deployment and scaling, resource management, data and application migration, testing and bench marking, standardization and interoperability, application and service management, and runtime and service discovery. This indicates that addressing one challenge may potentially impact the other areas, necessitating a holistic approach to tackle these issues. What makes these challenges unique to containerized multi-cloud setups is the dual complexity arising from container portability and cloud heterogeneity. Containers, while offering mobility and reproducibility, must now adapt to diverse runtime configurations, networking policies, and resource scheduling strategies across cloud providers. The multi-dimensionality of the challenges implies the need for an integrated approach that considers the interdependencies among these dimensions. It emphasizes the need for solutions that cater to this complexity rather than address individual challenges in isolation. Implications: For researchers, these findings underscore the need to develop multi-faceted, holistic solutions that can address the various interconnected challenges in managing multi-cloud environment. For practitioners, this implies the importance of strategic planning and implementing solutions that address the multiple facets of multi-cloud management while considering their potential impact on one another.
@@
Emergence of AI/ML Solutions: The framework highlights the increasing utilization of AI and ML in providing intelligent solutions to manage containerized applications in multi-cloud environment. This suggests the growing maturity and applicability of these technologies in complex scenarios, a significant step towards achieving efficient multi-cloud management. Previous studies (e.g., [164], [165], [166]) have recognized the potential of AI and ML in resolving complex cloud management issues, which is in line with the findings from the framework. This further validates the significant role AI and ML have started to play in this domain. The reviewed studies specifically show how AI/ML techniques help in real-time placement of containers across heterogeneous clouds, optimizing latency and cost under dynamic workload and policy constraints—issues that are intensified by the distributed and federated nature of multi-cloud environments. The increasing adoption of AI/ML in multi-cloud environment signifies an essential evolution in the field. This stresses the need for further research and development in these technologies to exploit their full potential in resolving multi-cloud management complexities. Implications: The growing use of AI/ML solutions implies that researchers need to focus on improving these technologies and adapting them to specific multi-cloud management challenges. For practitioners, this points towards the increasing necessity to invest in AI/ML capabilities and integrate them into their multi-cloud management strategies.
@@
Importance of Standardization and Interoperability: The framework underscores the crucial role of standardization and interoperability in multi-cloud environment, facilitating seamless integration and management across different cloud platforms. Recent literature validates the importance of standardization and interoperability in multi-cloud environment, emphasizing their role in enhancing the efficiency of cloud services [167]. This agrees with our results, further confirming the critical need for these factors in the domain. The need for standardization and interoperability reiterates the necessity for a unified framework or protocol across different cloud environments. In containerized deployments, interoperability challenges become more pronounced due to variations in networking stacks, security policies, and orchestration tools between providers. Standardization at the container runtime level (e.g., Docker images, Kubernetes manifests) and orchestration abstraction (e.g., TOSCA, Helm charts) plays a vital role in enabling automation workflows that span diverse cloud ecosystems. It suggests that attention to these aspects can lead to more seamless, efficient, and secure management of multi-cloud services. Implications: For researchers, the results stress the need to investigate methods to achieve better standardization and interoperability in multi-cloud environment. For practitioners, the findings suggest the need to adopt standards and focus on interoperability while selecting and implementing cloud services, ensuring more efficient management across different platforms.
4.5 Security Challenges and Solutions Framework (RQ4 - RQ5)
@@
Emphasize on Container-Specific Security Mechanisms: Security solutions specifically designed for containerized applications are critical to managing threats in multi-cloud environment. The selected studies suggest a variety of methods, including secure container orchestration, container isolation, and advanced deployment strategies that make use of deep reinforcement learning. This aligns with the existing literature (e.g., [168], [169] [170]) which highlights the unique security considerations brought forth by containerization. Containers, being lighter than traditional virtual machines and offering process-level isolation, have revolutionized the way applications are packaged and run. However, they also introduce new security challenges that need to be addressed by using container-specific security mechanisms. Containerization, while a powerful tool for application deployment, necessitates its own suite of security strategies. Implementing these will be key to leveraging the benefits of containerization in a secure manner. Implementing these will be key to using the benefits of containerization in a secure manner. In multi-cloud settings, these challenges are compounded by the need to secure container interactions across diverse platforms, enforce consistent policies under different providers, and protect workloads that dynamically move between clouds. Implications: For researchers, this highlights an area of potential study: the development of advanced container-specific security strategies, including those that use advanced techniques such as machine learning. For practitioners, this suggests the need to be well-versed in the specifics of container security. Leveraging features of containerization platforms, such as Docker’s isolation features, or incorporating tools specifically designed for container security will be crucial.
@@
Implementing Multi-Layered Security and Compliance Measures: Implementing multi-layered security measures that encompass data protection, access control, and communication security, along with compliance with legal and regulatory requirements, is of utmost importance. Solutions range from data encryption, role-based access control, secure container orchestration, adherence to regulations like GDPR [171]. This is consistent with current cybersecurity frameworks, such as the NIST cybersecurity framework [172], which emphasize a multi-layered approach to security. Furthermore, several studies point to the criticality of data security, user access control, and secure communication in cloud environments [173] [160]. In a multi-cloud, containerized environment, it is not enough to secure data; access and communication channels must also be secured, and all actions must comply with legal and regulatory requirements. In multi-cloud containerized setups, multi-layered security must also accommodate interoperability between CSPs, dynamic access management across domains, and distributed compliance enforcement for container workflows. Implications: For practitioners is to design and implement a comprehensive security strategy that encompasses multiple layers of protection. Compliance with legal and industry standards should be an integral part of this strategy. For researchers, this underlines the need for studies that address the integration of various layers of security and the development of comprehensive security frameworks.
@@
Ensuring Security through Standardization and Interoperability: Ensuring security through standardization and interoperability is highlighted as a critical step. Secure and standardized interfaces, such as OpenStack APIs, can be used to enhance security and performance across multiple cloud providers. This finding is aligned with existing literature, which emphasizes the role of standardization in achieving security and interoperability in multi-cloud environment [174] [175]. By implementing standardized interfaces and practices, organizations can ensure a level of interoperability and security across cloud environments. For containerized applications, this also means ensuring consistent security postures across orchestrators, managing identity federation between providers, and enabling trust in federated environments that dynamically provision containers at runtime. Implications: For practitioners, this means choosing platforms and tools that support standardized interfaces and practices, or working to implement those standards within their own environments. Researchers, on the other hand, could focus on the development of new standards or the improvement of existing ones to better address security challenges in multi-cloud, containerized environments.
4.6 Deployment Challenges and Solutions Framework (RQ4 - RQ5)